Find the magnitude and direction of the resultant
Answer #1

here,

writing each force according to their component in respective axis

Fx = Fa + Fb*cos60 = 5 + 10*0.5
Fx = 10 lb

Fy = Fbsin60 - Fc = 10*0.86 - 6
Fy = 2.6 lb

R = sqrt(Fx^2 + Fy^2)

R = sqrt(10^2 + 2.6^2)

R = 10.33 lb

Resultant of three vector is 10.33 lb

TanA = Fy/Fx
A = arcTan(2.6/10)

A = 14.57 degrees

Resultant Vector makes 14.57 degrees with positve X axis